OpenStudy (anonymous):

A set of water waves travels at 10.0 m/s, and 4.0 waves pass you in 2.0 s What is the wavelength of the waves? 5.0 m , 4.0 m , 0.25 m , or 0.2 m

OpenStudy (michele_laino):

please wait a moment

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

I'm pondering...

OpenStudy (anonymous):

ok!!

OpenStudy (michele_laino):

If I call with \lambda the wavelength of each wave, then we can write: \[\frac{{8\lambda }}{2} = 10\] from which I find \lambda=2.5 meters which is not an option of yours

OpenStudy (anonymous):

yes:/ i am not sure...but if i apply that equation, it should be like so? i got 10*2 =20/4=5? would it be 5 m is the solution?:

OpenStudy (anonymous):

Yes 5.0 m is the answer

OpenStudy (anonymous):

Did your question change from 8.0 waves to 4.0 waves?

OpenStudy (michele_laino):

if the answer is 5 meters then 8 waves are long 5*8 = 40 meters, and 40/2 = 20 m/second

OpenStudy (anonymous):

ooh okay! and so whenever we get these kinds of problems, we use that equation? oh yes, i put 8 earlier by accident, but i changed it to 4 :/ did it not go through on your side? :/

OpenStudy (anonymous):

Yes I noticed it changed so the answer is 5.0 m

OpenStudy (anonymous):

it shows as 4.0 on my side currently... what about on your side? :/

OpenStudy (anonymous):

oh okay! yay! sorry about that! :P OS is kind of wonky right now lol :P thank you1!

OpenStudy (anonymous):

No problem

OpenStudy (michele_laino):

maybe there is a typo in your options?